bella2025:

Solve the system by substitution: 3x + 2y = 8 x = 4 - 2y

Tranquility:

Substitute the second equation into the first one 3x + 2y = 8 x = 4 - 2y The second equation is x equals 4 - 2y So to solve it, replace 'x' in the first equation with 4 - 2y You then get 3(4-2y) + 2y = 8 Now you have an equation with just one variable which is y and you can find out the value of y. Can you simplify that equation and solve for y? First distribute the 3 inside the parenthesis 3(4-2y) + 2y = 8

bella2025:

12-6y+2y=8 12-4y=8

Tranquility:

Solve for y. Get y all by itself one side

Tranquility:

12 - 4y = 8 You can 4y on both sides and then subtract 8 on both sides

bella2025:

@tranquility wrote:
12 - 4y = 8 You can 4y on both sides and then subtract 8 on both sides
-4y=-4 1

Tranquility:

@bella2025 wrote:
-4y=-4 1
That is correct. y = 1 Since you know y = 1, you can plug this into either equation and now solve for x x = 4 - 2y If y = 1, then what is x = ??

bella2025:

@tranquility wrote:
@bella2025 wrote:
-4y=-4 1
That is correct. y = 1 Since you know y = 1, you can plug this into either equation and now solve for x x = 4 - 2y If y = 1, then what is x = ??
hang on

Tranquility:

alright

bella2025:

@tranquility wrote:
alright
I dont know

Tranquility:

x = 4 -2y We found out y = 1 Replace y with 1 and solve for x x = 4 - 2*(1)

bella2025:

4=-2y 0.5

Tranquility:

... no

Tranquility:

@tranquility wrote:
x = 4 -2y We found out y = 1 Replace y with 1 and solve for x x = 4 - 2*(1)
what does x equal?

bella2025:

3x - 2y = 8 3x-2(1)=8 3x-2=8 3x=10 3.33

Tranquility:

You wrote the equation incorrectly. It's 3x PLUS 2y = 8 Not 3x minus 2y = 8

bella2025:

3x + 2y = 8 3x+2(1)=8 3x+2=8 3x=6 x=2

axie:

Do you still need help with this?

Tranquility:

@bella2025 wrote:
3x + 2y = 8 3x+2(1)=8 3x+2=8 3x=6 x=2
That is correct